OpenStudy (anonymous):

e^(ln(x+5))=-6+2x solve for x CHECK ANSWER x=11?

OpenStudy (anonymous):

x+5=-6+2x 11=x gj

OpenStudy (anonymous):

sweet thanks mate

OpenStudy (anonymous):

np